# Rigidity of Gradient Shrinking Ricci Solitons with Bach-like Flatness and Related Variational Formulas

### v2

@@ 7 The Vanishing U-tensor
In order to extend the $(\alpha,\beta)$-cone in Theorem 1.2, we need to consider the implications of $\mathrm{U}_{ij}=0$. We attempt the same style of argument as we did with $\mathrm{V}_{ij}=0$. Define a weighted volume element of the form
@@
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1), normalized so that $\mathrm{R}+|\nabla f|^{2}=f$. Let $r$ be a regular value of $f$ and define the sub-level set $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. If $c>0$ is a constant and $\Delta\mathrm{R}=0$, then we have the following:
@@
    $\int_{\Omega_{r}}f\,\langle\nabla R,\nabla f\rangle\,\mathrm{d}V_{c}=0$.
@@
$=c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\,\mathrm{d}V.$
Next, we have:
@@
where we used 2.2(3). And the last critical identity is:
@@
$=0.$
We first state a quick consequence of Bochner’s formula, and then we will proceed with the main integral identity.
@@
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1), normalized so that $\mathrm{R}+|\nabla f|^{2}=f$. Then if $\Delta\mathrm{R}=0$:
$2|\mathrm{Rc}|^{2}=\mathrm{R}+\langle\nabla\mathrm{R},\nabla f\rangle$
@@
$\frac{1}{2}(2-\mathrm{R})$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+|\frac{1}{2}g-\mathrm{Rc}|^{2}$
@@
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Assume that $\Delta\mathrm{R}\equiv 0$, $r>0$ is a regular value of $f$, and $\Omega_{r}=\{f\leq r\}$. Then
@@
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}\,|\nabla f|^{2}-\mathrm{R}\,|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
@@
$U_{ij}=\mathrm{R}_{ij}+\nabla_{k}\mathrm{R}_{ij}\nabla_{k}f-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}g_{ij}+\frac{\mathrm{R}^{2}}{4}g_{ij},$
@@
    $\mathrm{\mathrm{Rc}}(\nabla f)=\frac{1}{2}\nabla\mathrm{R}\implies\mathrm{\mathrm{Rc}}(\nabla f,\nabla f)=\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$,
@@
The second term in our expression for $U$ is simplified by looking at the following divergence:
@@
$=\int_{\Omega_{r}}\mathrm{div}(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{k}f\,\mathrm{e}^{-cf})\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
@@
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{d}V_{c}.$
Now we can transform the integral using our Bochner formula from Lemma 7.2 as well as
@@
does not yield any obvious rigidity. The $\mathrm{R}^{2}|\nabla f|^{2}$ term reminds us of our work on the $\mathrm{V}$-tensor, but we must first reformulate our integral identity for $\mathrm{V}_{ij}$ for the altered-weighted volume element.
@@
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Let $r>0$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. Then if $\Delta\mathrm{R}=0$ and $c>0$, we have
$\int_{\Omega_{r}}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
and
@@
We will once again integrate the terms of $\mathrm{V}(\nabla f,\nabla f)$ separately. First, we compute the following divergence:
@@
Hence, the integral of $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}$ is:
@@
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
@@
$=\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\mathrm{e}^{-cr}\Big(\frac{c-1}{c}+1-\frac{1}{2c}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V-\Big(\frac{c-1}{c}+\frac{1}{2}-\frac{1}{2c}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
@@
Now taking the limit $r\rightarrow\infty$:
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V=-\int_{M}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{e}^{-cf}\,\mathrm{d}V.$
8 Proof of Theorem 1.1(ii)
Assume $\mathrm{U}_{ij}\equiv 0$. By Lemma 2.1, $\mathrm{tr}_{g}(\mathrm{U})=-\Delta\mathrm{R}$, hence $\Delta R\equiv 0$. Fix a regular value $r\in(0,1)$ of $f$ and write $\Omega_{r}=\{x\in M:f(x)\leq r\}$. For any $c>0$, Lemma 7.3 gives
$0=\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\!\Big((\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}\!|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
Multiplying by $4c$ and rearranging yields the basic identity
$c\int_{\Omega_{r}}(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\big(\mathrm{e}^{-cr}-\mathrm{e}^{-cf}\big)\,\,\mathrm{d}V.$
On the fixed bounded set $\Omega_{r}$ we have $0\leq f\leq r<1$, so all integrands are smooth and dominated by an $L^{1}$ function independent of $c$. Hence we may differentiate under the integral sign. Differentiating once gives
$\int_{\Omega_{r}}(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V-c\int_{\Omega_{r}}f\,(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(f\,\mathrm{e}^{-cf}-r\,\mathrm{e}^{-cr}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
Differentiating again yields
$\int_{\Omega_{r}}\big(cf^{2}-2f\big)(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(r^{2}\,\mathrm{e}^{-cr}-f^{2}\,\mathrm{e}^{-cf}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
An easy induction shows that for each integer $n\geq 0$,
$\int_{\Omega_{r}}\big(nf^{\,n-1}-cf^{\,n}\big)\,(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(r^{\,n}\,\mathrm{e}^{-cr}-f^{\,n}\,\mathrm{e}^{-cf}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$ (8)
Setting $c=0$ in (8) and summing over $n\geq 0$, we use the geometric series sum formulas:
$\sum_{n=0}^{\infty}nf^{\,n-1}=\frac{1}{(1-f)^{2}},\qquad\sum_{n=0}^{\infty}f^{\,n}=\frac{1}{1-f},\qquad\sum_{n=0}^{\infty}r^{\,n}=\frac{1}{1-r}.$
This leads to the identity
$\int_{\Omega_{r}}\Bigg(\frac{(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}}{(1-f)^{2}}+\frac{f-r}{(1-f)(1-r)}\,|\nabla\mathrm{R}|^{2}\Bigg)\,\,\mathrm{d}V=0.$ (9)
On $\Omega_{r}$ we have $0\leq\mathrm{R}\leq f<1$, hence $\mathrm{R}^{2}-\mathrm{R}\leq 0$ and $f-r\leq 0$, while (1f)>0(1-f)>0 and (1r)>0(1-r)>0. Therefore both terms in the integrand of (9) are non-positive. Since the integral vanishes, both terms vanish identically on $\Omega_{r}$:
[…]
In particular $\nabla\mathrm{R}\equiv 0$ on $\Omega_{r}$. Together with $\Delta\mathrm{R}\equiv 0$, unique continuation implies $\mathrm{R}$ is constant on all of $M$. If $\mathrm{R}=0$ at some point then $\mathrm{R}\equiv 0$ everywhere and the soliton is isometric to the Gaussian soliton on $\mathbb{R}^{4}$. Otherwise $\mathrm{R}>0$ and the soliton is Einstein.
9 Proof of Theorem 1.2 Outside the Cone
Assume $3\beta\neq\alpha$, then $\mathrm{tr}_{g}\mathfrak{B}_{ij}=(3\beta-\alpha)\Delta\mathrm{R}=0$, hence $\Delta\mathrm{R}\equiv 0$. For any $c>0$ and a fixed regular value $r\in(0,1)$ for $f$, let $\Omega_{r}=\{f\leq r\}.$ We express the Bach-like tensor as
$\mathfrak{B}_{ij}=6\beta\,\mathrm{B}_{ij}+(\alpha-3\beta)\,\mathrm{U}_{ij}$
One can easily check that for the volume element $\mathrm{d}V_{c}$, the integration of the Bach tensor yields:
$\int_{\Omega_{r}}\mathrm{B}(\nabla f,\nabla f)\,\,\mathrm{d}V_{c}=-\frac{c}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V_{c}=\mathrm{F}(c).$
So differentiating past the integral symbol yields
$\frac{d^{n}}{dc^{n}}\Big|_{c=0}\mathrm{F}(c)=-\frac{1}{2}\int_{\Omega_{r}}nf^{n-1}\,|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V.$
We will need to multiply both sides of the Bach-like equation by $c$ to proceed.
$\frac{d^{n}}{dc^{n}}\Big(c\,\mathrm{F}(c)\Big)=n\,\mathrm{F}^{(n-1)}(c)+c\,\mathrm{F}^{(n)}(c),$
hence
$\frac{d^{n}}{dc^{n}}\Big|_{c=0}=n\,\mathrm{F}^{(n-1)}(0)=-\frac{1}{2}\int_{\Omega_{r}}n(n-1)\,f^{n-2}\,|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V$
Summing over $n$ on the right-hand side, we obtain
$\int_{\Omega_{r}}\frac{|\mathrm{D}_{ijk}|^{2}}{(1-f)^{3}}\,\mathrm{d}V.$ (10)
Now assuming $\mathfrak{B}_{ij}\equiv 0$, and multiplying both sides by $c$:
$0=6\beta\,c\int_{\Omega_{r}}\mathrm{B}(\nabla f,\nabla f)\,\mathrm{d}V_{c}+(\alpha-3\beta)\,c\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c},$
hence
$0=\frac{-6\beta c^{2}}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V_{c}+(\alpha-3\beta)\,c\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}.$
Now differentiating with respect to $c$, evaluating each derivative at $c=0$ and summing over $n$ we finally obtain
$\boxed{-6\beta\int_{\Omega_{r}}\frac{|\mathrm{D}_{ijk}|^{2}}{(1-f)^{3}}\,\mathrm{d}V+\frac{1}{4}(\alpha-3\beta)\int_{\Omega_{r}}\Bigg(\frac{(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}}{(1-f)^{2}}+\frac{f-r}{(1-f)(1-r)}\,|\nabla\mathrm{R}|^{2}\Bigg)\,\,\mathrm{d}V=0.}$
Now as $0<r<1$, both terms in the second integrand are negative as in (9). If $\beta>0$ and $\alpha-3\beta>0$, or if $\beta<0$ and $\alpha-3\beta<0$, then
$|\mathrm{D}_{ijk}|^{2}\equiv 0,\quad\mathrm{R}^{2}-\mathrm{R}\equiv 0,\quad\mathrm{and}\quad|\nabla\mathrm{R}|^{2}\equiv 0\quad\mathrm{on}\,\,\Omega_{r}.$
Hence, $M$ must be Einstein or isometric to the Gaussian shrinking soliton. These ranges for $\alpha$ and $\beta$ give us the desired result for the region outside the cone $\mathcal{C}$, hence for all $(\alpha,\beta)$ aside from $(0,0)$ and the Bach-line.

### v3

@@ 7 The Vanishing U-tensor
The analysis we used for the vanishing $V$-tensor doesn’t completely carry over to $\mathrm{U}=0$, but will allow us to prove Theorem 1.2. Define a weighted volume element of the form
@@
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Let $r$ be a regular value of $f$ and define the sub-level set $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. If $c>0$ is a constant and $\Delta\mathrm{R}=0$, then we have the following:
@@
    $\int_{\Omega_{r}}f\,\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=0$.
@@
$=c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\,\mathrm{d}V,$
showing (1). Next, we have:
@@
showing (2). And the last identity is:
@@
$=0,$
showing (3). ∎
We first state a quick consequence of Bochner’s formula.
@@
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). If $\Delta\mathrm{R}=0$, then:
$2|\mathrm{Rc}|^{2}=\mathrm{R}+\langle\nabla\mathrm{R},\nabla f\rangle.$
@@
$\frac{1}{2}(2-\mathrm{R})$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+\left|\frac{1}{2}g-\mathrm{Rc}\right|^{2}$
@@
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Assume that $\Delta\mathrm{R}\equiv 0$, $r>0$ is a regular value of $f$, and $\Omega_{r}=\{f\leq r\}$. Then
@@
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}\,|\nabla f|^{2}-\mathrm{R}\,|\nabla f|^{2}+\tfrac{1}{c}|\nabla\mathrm{R}|^{2}\big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
@@
$\mathrm{U}_{ij}=\mathrm{R}_{ij}+\nabla_{k}\mathrm{R}_{ij}\nabla_{k}f-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}\,g_{ij}+\frac{\mathrm{R}^{2}}{4}g_{ij},$
@@
    $\mathrm{Rc}(\nabla f)=\frac{1}{2}\nabla\mathrm{R}\implies\mathrm{Rc}(\nabla f,\nabla f)=\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$,
@@
The second term in our expression for $\mathrm{U}$ is simplified by looking at the following divergence:
@@
$=\int_{\Omega_{r}}\mathrm{div}\left(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{k}f\,\mathrm{e}^{-cf}\right)\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
@@
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{d}V_{c}.$
Now transforming using the Bochner formula from Lemma 7.2:
@@
does not yield obvious rigidity. For comparison, we state the analogous identity for $\mathrm{V}(\nabla f,\nabla f)$ under $\mathrm{d}V_{c}$. The contrast with Lemma 7.3 is instructive: for $c\geq 1$ the integrand for $\mathrm{V}$ acquires a nonnegative term, while the integrand for $\mathrm{U}$ remains sign-indefinite for any $c>0$. This further highlights the subtleties in the $\mathrm{U}=0$ case.
@@
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Let $r>0$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. If $\Delta\mathrm{R}=0$ and $c>0$, then
$\int_{\Omega_{r}}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V,$
and in particular
@@
We compute the following divergence:
@@
Hence the integral of $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}$ is:
@@
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S$
@@
$=\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
@@
Taking the limit $r\rightarrow\infty$ gives the stated global identity. ∎
8 Proof of Theorem 1.2
Proof.
Assume that $\mathrm{U}_{ij}=0$. By Lemma 2.1, $\Delta\mathrm{R}=0$. Let $p$ be the point where $f$ attains its minimum value $f_{\min}$. Then $\nabla f(p)=0$, so by (2), $\mathrm{R}(p)=f_{\min}$. Using $\mathrm{U}=0$ at $p$ with $\Delta\mathrm{R}=0$:
$\mathrm{Rc}(p)\big(1-\mathrm{R}(p)\big)+\tfrac{1}{4}\mathrm{R}(p)\big(\mathrm{R}(p)-1\big)g(p)=0,$
which factors as
$\big(1-\mathrm{R}(p)\big)\!\left(\mathrm{Rc}(p)-\tfrac{1}{4}\mathrm{R}(p)\,g(p)\right)=0.$
If $\mathrm{R}(p)\neq 1$ then $\mathrm{Rc}(p)=\tfrac{1}{4}\mathrm{R}(p)g(p)$, which gives
$|\mathrm{Rc}(p)|^{2}=\tfrac{1}{4}\mathrm{R}^{2}(p).$
Since $\nabla f(p)=0$, Lemma 7.2 gives $2|\mathrm{Rc}(p)|^{2}=\mathrm{R}(p)$. Substituting:
$\tfrac{1}{2}\mathrm{R}^{2}(p)=\mathrm{R}(p)\implies\mathrm{R}(p)\big(\mathrm{R}(p)-2\big)=0.$
Hence $\mathrm{R}(p)\in\{0,2\}$. Together with the $\mathrm{R}(p)=1$ case, we have $f_{\min}=\mathrm{R}(p)\in\{0,1,2\}$.
Case $f_{\min}=0$. We have $\mathrm{R}(p)=0$, and the manifold is isometric to the Gaussian soliton by [8].
Case $f_{\min}=2$. Using the soliton equation,
$\mathrm{R}^{2}-2|\mathrm{Rc}|^{2}=\mathrm{R}^{2}-2\left|\tfrac{1}{2}g-\nabla^{2}f\right|^{2}=\mathrm{R}^{2}-2\mathrm{R}+2-2|\nabla^{2}f|^{2}.$
At the minimum $p$ with $f_{\min}=2$, we have $\Delta f(p)=2-\mathrm{R}(p)=0$. Since $p$ is a minimum, $\nabla^{2}f(p)\geq 0$, and $\Delta f(p)=0$ forces $\nabla^{2}f(p)=0$. Thus:
$\mathrm{R}^{2}(p)-2\mathrm{R}(p)+2-2|\nabla^{2}f(p)|^{2}=4-4+2-0=2>0.$
By continuity, there exist $r>2$ and $\epsilon>0$ such that $\mathrm{R}^{2}-2\mathrm{R}+2-2|\nabla^{2}f|^{2}\geq\epsilon>0$ on $\Omega_{r}$. Therefore, under $\mathrm{U}\equiv 0$,
$\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V=0$
implies $|\nabla f|^{2}\equiv 0$ on $\Omega_{r}$, hence $\mathrm{R}=f-|\nabla f|^{2}=2$ on $\Omega_{r}$. By unique continuation for harmonic functions, $\mathrm{R}=2$ on all of $M$, and the soliton is Einstein by [8].
For the converse, if $M$ is Einstein then $\mathrm{B}\equiv 0$ (as $\mathrm{B}$ vanishes on Einstein manifolds [4]) and $\mathrm{V}\equiv 0$ by Theorem 1.1, so $\mathrm{U}=2\mathrm{B}-\tfrac{1}{3}\mathrm{V}\equiv 0$. If $M$ is Gaussian then all curvature vanishes and $\mathrm{U}\equiv 0$.
Case $f_{\min}=1$. At a minimum point $p$ the $\mathrm{U}$-tensor vanishes independently of any global assumption, and $\mathrm{V}$ cannot vanish. Coupled with the sign issue for the $\mathrm{U}$-vanishing integral identity, this case remains open; see Conjecture 1.3. ∎